12. A professor computes the mean difference in exam scores before and after a lecture as 9.0 ± 14.5 (MD ± SD), and this difference was significant. What is the effect size for this result using estimated Cohen's d? A) d=0.62 (medium effect) B) d=1.61 (medium effect) C) d=0.62 (large effect) D) d=1.61 (large effect) E) d=0.62 (large effect)
Answer #1

since we know that Cohen's D =MD/SD =9/14.5=0.62

option A is correct

A) d=0.62 (medium effect)

(since it is between 0.50 and 0.80)